# Black-box Adversarial Prompting

### v1

@@ 4.3 Iterative Refinement
In subsequent iterations, the generator synthesizes new adversarial prompt suffixes conditioned on the top-$k$ exemplars from the previous round. After several iterations, the prompt candidate $\mathcal{P}_{adv}^{*}$ that yields the lowest average response length on $\mathcal{D}^{\prime}$ while maintaining competitive accuracy is selected as the final adversarial prompt for deployment.
@@ 5 Experiments
AdvPrompt 96.1 440 28.1% 95.2 2176 49.5% 96.9 4019 57.9% 70.0 8659 76.1% 89.6 63.0%
Qwen3-32B
AdvPrompt 96.2 435 27.2% 95.3 2295 51.8% 96.2 4125 60.2% 73.3 8700 79.8% 90.3 65.4%
Table 1: Experimental results of AdvPrompt across various types of reasoning models. We report accuracy (Acc.), average token usage (Tok.), and compression ratio for comparison. Best efficiency–performance trade-offs are highlighted in bold. We include DEER, a white-box inference intervention method, for reference purposes only.
@@ 5.1 Experimental Setup
We evaluate our method on both open-source LRMs and closed-source APIs. The open-source models include the DeepSeek-R1-Distill-Qwen model series DeepSeek-AI et al. (2025) and the Qwen3 series Yang et al. (2025a). For closed-source APIs, we assess Claude-3.7-Sonnet-Thinking [^5] We use the Claude-3.7-Sonnet-20250219-Thinking version for experiments. Anthropic (2024) and Gemini-2.5-Pro-Thinking  Comanici et al. (2025). Evaluation is conducted across four widely used mathematical reasoning benchmarks: GSM8K (Cobbe et al., 2021), MATH Hendrycks et al. (2021), AMC 2023, and AIME 2024. Regarding the MATH dataset, due to the computation cost, we assess our method on a subset, MATH-500, which is identical to the test set used in Lightman et al. (2024). We construct our PDSet by randomly sampling 100 instances from the math split of PRM800K, the training data used in the same work.
@@
Inference for our proposed method and all baselines is conducted using the vLLM [^6] https://github.com/vllm-project/vllm package. We maintain a sampling temperature of $0.6$, a top-p value of $0.95$, and permit a maximum of 16,384 tokens to be generated. The number of samplings during evaluation depends on the dataset size: 3 samples per question for GSM8K and MATH-500, and 8 samples for AMC 2023 and AIME 2024. For AdvPrompt, we utilize the top-$5$ candidates as exemplars for each next iteration and set the tolerance threshold $\tau$ to $1.0$; the number of refinement iterations is set to $3$. We include more implementation details in Appendix B.1.
@@
[…]
@@ 5.2 Results on Open-source LRMs
Contrary to the common belief that mitigating overthinking in LRMs requires specialized training or inference-time interventions (e.g., early exiting), we demonstrate that leveraging the instruction-following capabilities of LRMs alone can substantially improve reasoning efficiency. As shown in Table 1, compared to existing prompting baselines that suffer from substantial performance degradation or limited efficiency, AdvPrompt achieves an average token reduction of $34\%$$37\%$ on the Qwen3 series across four benchmarks, and up to $22\%$ on the DeepSeek-R1-Distill-Qwen-14B model while preserving the reasoning performance of LRMs. [^7] See Appendix B.3 for additional results on other model scales of DeepSeek-R1-Distill-Qwen and Qwen3 series. The experimental results also illustrate the superiority of AdvPrompt over DEER, a white-box inference intervention approach, with a relative improvement of $6\%$$24\%$ in compression ratio.
Consistent with prior studies Yang et al. (2025c); Liu et al. (2025), we observe that AdvPrompt is particularly effective on simpler reasoning tasks. Notably, it achieves nearly a $3\times$ reduction in response length on GSM8K and a $2\times$ reduction on MATH-500. These results align with the broader objective of efficient reasoning Sui et al. (2025), which advocates allocating greater computational resources to complex problems while encouraging concise reasoning for simpler ones.
@@ 5.3 Results on Closed-source APIs
Figure 3 presents the experimental results of applying AdvPrompt to two widely used commercial APIs. Due to budgetary constraints, the evaluation is primarily conducted on the MATH-500 benchmark. As shown in the figure, AdvPrompt reduces average token usage by $35\%$ on Claude-3.7-Sonnet-Thinking and by $47\%$ on Gemini-2.5-Pro-Thinking. Importantly, AdvPrompt preserves the original reasoning performance of these high-performing APIs, with less than $3\%$ accuracy drop on Claude and nearly NO degradation on Gemini. These results highlight the effectiveness of AdvPrompt as a general-purpose, black-box solution for efficient reasoning with closed-source reasoning models.
Figure 3: Experimental results of AdvPrompt using closed-source APIs on MATH-500. It effectively reduces token usage by $35\%$$47\%$ while maintaining comparable reasoning performance.
@@ 6 Analysis
In this section, we provide a comprehensive analysis of AdvPrompt, covering the effectiveness of diverse adversarial perspectives (§6.1), the generalizability of adversarial prompts across different model scales and families (§6.2), model-specific sensitivity to AdvPrompt (§6.3), and the impact of iterative refinement (§6.4).
Figure 4: Compression ratios (%) of top-55 prompt candidates across perspectives. The Evidence perspective is most effective on Qwen3-8B, while RolePlay candidates perform best on Deepseek-R1-Distill-Qwen-14B.
@@ 6.1 Effectiveness of Different Perspectives
Figure 4 illustrates the compression ratios of the top-$5$ prompt candidates generated from each adversarial perspective, averaged across four benchmarks. Additional experimental results for the Qwen3 and DeepSeek-R1-Distill-Qwen model series are provided in Appendix B.4. The results show that prompts based on the Evidence-based persuasion technique consistently achieve superior compression within the Qwen3 series. Specifically, the top-$4$ Evidence-based candidates for Qwen3-8B yield average compression ratios ranging from $66\%$ to $70\%$, marking the best performance across all evaluated perspectives. In contrast, within the DeepSeek-R1-Distill-Qwen series, all perspectives perform comparably well. For instance, on DeepSeek-R1-Distill-Qwen-14B, the top-$1$ candidates from four perspectives—excluding that from Emotion perspective—yield similar compression ratios between $78\%$ to $80\%$. Across both model families, prompts derived from the Threat perspective consistently underperform relative to others, with average token reductions ranging from $4.2\%$ to $25.8\%$. Representative examples of high-performing prompts from each perspective are provided in Appendix A.3.
@@ 6.2 Generalizability Analysis
Figure 5 shows the generalizability of AdvPrompt across different model scales within the Qwen3 family. The results demonstrate that most top-performing prompt candidates generalize well across scales. For instance, Evidence-I [Uncaptioned image] consistently ranks among the top-$5$ candidates across all Qwen3 variants, while Evidence-III [Uncaptioned image] ranks within the top-$2$ candidates for the 4B, 8B, and 14B models. [^8] Detailed prompts corresponding to specific indices are provided in Table 2. This strong intra-series generalizability is likely attributable to the shared pre-training corpora, architectural configurations, and training procedures across model scales within the same series.
@@
We further assess the generalizability of AdvPrompt across different model families. As reported in Appendix A.3, the Qwen3 and DeepSeek-R1-Distill-Qwen series share several top-performing prompt candidates, such as Evidence-II and RolePlay-III. This overlap suggests that certain adversarial prompt strategies possess general utility, highlighting the promise of AdvPrompt as a model-agnostic approach for efficient reasoning in black-box LRMs.
Figure 6: Sensitivity of different models to AdvPrompt. Solid lines represent mean results; black diamonds indicate outliers. Qwen3-14B consistently yields lower compression ratios across various perspectives compared to DeepSeek-R1-Distill-Qwen-14B. Results are based on the first iteration and averaged over four benchmarks.
@@ 6.3 Sensitivity Analysis
Figure 6 compares the sensitivity of the primary evaluated model series to AdvPrompt. The results indicate that the Qwen3 series is more responsive to adversarial prompting than the DeepSeek-R1-Distill-Qwen series. In particular, Qwen3-14B consistently achieves lower compression ratios across various adversarial perspectives compared to DeepSeek-R1-Distill-Qwen-14B, with average improvements ranging from $4\%$ to $12\%$. This trend is consistent with the findings in Table 1, where Qwen3-14B attains an average compression ratio of $63\%$, whereas DeepSeek-R1-Distill-Qwen-14B achieves $78\%$.
@@ 6.4 Impact of Refinement Iteration
We further examine the impact of iterative refinement on the performance of AdvPrompt. As illustrated in Figure 7, successive refinement rounds improve the compression ratio of the best-performing prompt candidates. Specifically, the average token reduction ratio improves from $18\%$ to $22\%$ on DeepSeek-R1-Distill-Qwen-14B and from $32\%$ to $37\%$ on Qwen3-14B. Importantly, these refinements do not compromise the original reasoning accuracy, indicating the robustness of AdvPrompt across multiple iterations. No further gains are observed beyond three iterations. Accordingly, we adopt three refinement rounds in our experiments.
@@ 7 Conclusion
This work presents AdvPrompt, an iterative refinement framework that generates high-quality adversarial prompts from diverse perspectives to elicit concise responses of LRMs without compromising their accuracy. Experiments on both LRMs and closed-source APIs validate its effectiveness in reducing token usage while preserving reasoning performance. Further analysis highlights the generalizability of AdvPrompt across different model scales and families. We hope this study offers new insights into efficient reasoning and underscores the potential of black-box adversarial prompting as a practical strategy for improving LRM efficiency.
@@
Due to computational constraints, we did not conduct experiments on larger LRMs such as Qwen3-235B-A22B. We believe that AdvPrompt could retain its effectiveness in enhancing the reasoning efficiency of such models. The open-source LRMs evaluated in this study are primarily from the Qwen3 and DeepSeek-R1-Distill-Qwen series. Future work will extend our investigation to a broader range of models, including the gpt-oss series OpenAI (2024). Nonetheless, the strong performance of AdvPrompt on closed-source APIs supports its generalizability across diverse model families. We plan to explore these aspects further, as we anticipate that AdvPrompt ’s potential can be more fully realized in these settings.
@@
As stated in Section 3, our work focuses on promoting concise and efficient reasoning in reasoning models, with a clear commitment to ensuring safe and ethical usage. While our method involves techniques that may bear surface resemblance to adversarial prompting, particularly in their potential to influence model behavior, it fundamentally diverges from conventional jailbreak-style adversarial attacks. Rather than eliciting harmful or undesired outputs, our approach is designed to enhance reasoning efficiency through principled black-box prompting. We fully support the development of adaptive safety mechanisms and advocate for transparent adversarial research to promote the secure and responsible evolution of LLM technologies.
@@
Appendix A Details of AdvPrompt
In this section, we provide details of our proposed AdvPrompt framework, including formal definitions of adversarial perspectives (§A.1), detailed instructions for prompt creation (§A.2), and representative examples of prompt candidates (§A.3).
@@ A.1 Definitions
We present detailed definitions of our adopted adversarial perspectives as follows:
  • Emotional Appeal. A persuasive technique that seeks to influence audiences by eliciting emotions, such as sadness, hope, or empathy, rather than relying exclusively on logic or factual evidence Petty et al. (2003).
  • Threat. A coercive form of persuasion that leverages explicit or implicit negative consequences, such as fear of punishment or loss, to compel behavioral or attitudinal compliance Johannesen and Larson (1989).
  • Evidence-based Persuasion. A rational appeal that utilizes credible sources, such as statistics, empirical findings, and expert testimony, to substantiate claims and influence beliefs, attitudes, or actions O’Keefe (2016).
  • Role-Playing. A strategy involving the enactment of scenarios or personas to foster empathy, encourage perspectives, and increase the likelihood of attitude or behavior change Chen et al. (2024a).
Due to computational constraints, this study investigates a limited set of adversarial perspectives. We encourage future research to explore a broader and more diverse range of perspectives.

### v2

@@ 4.3 Iterative Refinement
In subsequent iterations, the generator synthesizes new persuasive prompt suffixes conditioned on the top-$k$ exemplars from the previous round. After several iterations, the prompt candidate $\mathcal{P}_{adv}^{*}$ that yields the lowest average response length on $\mathcal{D}^{\prime}$ while maintaining competitive accuracy is selected as the final persuasive prompt for deployment.
@@ 5 Experiments
DeepSeek-R1-Distill-LLaMA-8B
Original 92.0 1715 100% 89.2 4102 100% 89.4 $60$ 100% 43.3 11015 100% 78.5 100%
NoThinking 80.0 266 15.5% 69.9 972 23.7% 58.1 1495 24.7% 15.8 4575 41.5% 56.0 31.9%
BeConcise 91.3 1482 86.4% 89.5 3753 91.5% 87.8 5660 93.4% 45.4 11236 102% 78.5 96.7%
ChainofDraft 90.2 1101 64.2% 89.2 3557 86.7% 88.1 5338 88.1% 44.6 11193 102% 78.0 92.6%
GEPA 89.8 1045 60.9% 88.6 3446 84.0% 88.4 5110 84.3% 46.8 10279 93.3% 78.4 86.8%
DEER 88.7 909 53.0% 86.2 3021 73.6% 86.9 4915 81.1% 44.0 10377 94.2% 76.5 84.0%
Whisper 91.4 624 36.4% 89.3 2823 68.8% 89.4 4722 77.9% 45.8 10210 92.7% 79.0 80.3%
@@
Whisper 96.1 440 28.1% 95.2 2176 49.5% 96.9 4019 57.9% 70.0 8659 76.1% 89.6 63.0%
Table 1: Experimental results of Whisper across various types of reasoning models. We report accuracy (Acc.), average token usage (Tok.), and compression ratio for comparison. Best efficiency–performance trade-offs are highlighted in bold. We include DEER, a white-box inference intervention method, for reference purposes only.
@@ 5.1 Experimental Setup
We evaluate our method on both open-source LRMs and closed-source APIs. The models include the DeepSeek-R1-Distill series deepseekr1 and Qwen3 series Qwen3. For commercial APIs, we assess Claude-3.7-Sonnet-Thinking [^5] We use the Claude-3.7-Sonnet-20250219-Thinking version for experiments. sonnet3_5 and Gemini-2.5-Pro-Thinking  gemini25. The evaluation is primarily conducted across four widely used mathematical reasoning benchmarks: GSM8K (Cobbe:2021gsm8k), MATH math, AMC 2023, and AIME 2024. Regarding the MATH dataset, due to computation cost, we assess our method on a subset, MATH-500, which is identical to the test set used in Lightman:2024verify. We construct our PDSet by randomly sampling 100 instances from the math split of PRM800K, the training data used in the same work.
@@
Inference for Whisper and all baselines is conducted using the vLLM [^6] https://github.com/vllm-project/vllm package. For Whisper, we utilize the top-$5$ candidates as exemplars for each next iteration and set the tolerance threshold $\tau$ to $1.0$; the number of refinement iterations is set to $3$. We include more implementation details in Appendix B.1.
@@
We compare our proposed Whisper against five baselines: 1) NoThinking prompts. LRMs are forced to bypass reasoning and directly output the answer. 2) Token-efficient prompts. Following Lee:2025Complexity, we append “Be concise.” to the initial instruction, encouraging brevity in LRM reasoning. 3) Budget control. We adopt Chain-of-Draft Xu:2025chainofdraft, which constrains LRMs to reason within a predefined length budget. 4) Prompt optimization. We include GEPA GEPA, an advanced reflective prompt optimization framework for comparison. The implementation is based on DSPy khattab2024dspy. 5) Inference intervention. We report results for DEER Yang:2025deer, a white-box inference intervention method, as a point of reference in our main results. Specifically, DEER elicits intermediate answers at potential reasoning transition points (e.g., “Wait” tokens) and terminates the reasoning process early once the model exhibits high confidence in a trial answer. Details of these baselines are provided in Appendix B.2.
@@ 5.2 Results on Open-source LRMs
As illustrated in Table 1, compared to existing simple prompting baselines (e.g., “Be concise.”) that suffer from substantial performance degradation or limited efficiency, Whisper achieves an average token reduction of $\bf{37\%}$ on Qwen3-14B across four benchmarks, and up to $\bf{22\%}$ on the DeepSeek-R1-Distill model, while maintaining comparable reasoning performance. [^7] See Appendix B.3 for additional results on other model scales of DeepSeek-R1-Distill and Qwen3 series. Additionally, Whisper consistently outperforms GEPA across various models, demonstrating the superiority of diverse persuasive perspectives over conventional prompt optimization. The experimental results also show that Whisper outperforms DEER, a white-box inference intervention approach, achieving an absolute improvement of up to $\bf{18\%}$ in compression ratio. Importantly, as a black-box prompting method, Whisper is orthogonal to existing white-box approaches, highlighting its potential for further improvement through complementary integration.
Consistent with prior studies Yang:2025deer; Liu:2025Laser, we observe that Whisper is particularly effective on simpler reasoning tasks. Notably, across all the evaluated models, Whisper achieves nearly a $\bf{3\times}$ reduction in response length on simple GSM8K questions and an $\bf{1.4\times}$$\bf{2\times}$ reduction on MATH-500. These results align with the broader objective of efficient reasoning Sui:2025survey, which advocates allocating greater computational resources to complex problems while encouraging concise reasoning for simpler ones.
@@ 5.3 Results on Closed-source APIs
Figure 3 presents the results of applying Whisper to two widely used commercial APIs. Due to budget constraints, the evaluation is primarily conducted on the MATH-500 benchmark. As shown in the figure, Whisper reduces average token usage by $\bf{46\%}$ on Claude-3.7-Sonnet-Thinking and by $\bf{50\%}$ on Gemini-2.5-Pro-Thinking. Importantly, Whisper preserves the original reasoning performance of both high-performing APIs. These results highlight the effectiveness of Whisper as a general-purpose, black-box solution for efficient reasoning with closed-source models.
Figure 3: Experimental results with commercial APIs on MATH-500. Whisper effectively achieves a $\bf{2\times}$ token reduction while maintaining comparable performance.
@@ 6 Analysis
This section provides a comprehensive analysis of Whisper, covering the effectiveness of diverse perspectives (§6.1), the generalizability of persuasive prompts across model scales, families, and data domains (§6.2), model sensitivity to Whisper (§6.3), and the impact of iterative refinement (§6.4).
Figure 4: Compression ratios (%) of top-55 prompt candidates across perspectives. The Evidence perspective is most effective on Qwen3-14B, while RolePlay candidates perform best on Deepseek-R1-Distill-Qwen-14B.
@@ 6.1 Effectiveness of Different Perspectives
Figure 4 illustrates the compression ratios of the top-$5$ prompt candidates generated from each persuasive perspective, averaged across four benchmarks. [^8] Additional results for other Qwen3 and DeepSeek-R1-Distill models are provided in Appendix B.4. The results show that prompts based on the Evidence-based persuasion consistently achieve superior compression within the Qwen3 series. Concretely, top-$4$ Evidence-based candidates for Qwen3-14B yield average compression ratios ranging from $63\%$ to $65\%$, marking the best performance across all perspectives. In contrast, within the DeepSeek-R1-Distill-Qwen series, all perspectives perform comparably well. For instance, on DeepSeek-R1-Distill-Qwen-14B, the top-$1$ candidates from four perspectives—excluding that from Emotion perspective—yield similar compression ratios between $78\%$ to $80\%$. Representative examples of high-performing prompts from each perspective are provided in Appendix A.3.
@@ 6.2 Generalizability Analysis
Figure 5 shows the generalizability of Whisper across different model scales within the Qwen3 family. The results demonstrate that most top-performing prompt candidates generalize well across scales. For instance, Evidence-I [Uncaptioned image] consistently ranks among the top-$5$ candidates across all Qwen3 variants, while Evidence-III [Uncaptioned image] ranks within the top-$2$ candidates for the 4B, 8B, and 14B models. [^9] For detailed prompts corresponding to specific indices, refer to Table 3 in Appendix A.3. This strong intra-series generalizability is likely attributable to the shared pre-training corpora, architectural configurations, and training procedures across model scales within the same series.
@@
We further assess the generalizability of Whisper across different model families. As shown in Appendix A.3, the Qwen3 and DeepSeek-R1-Distill-Qwen series share several top-performing prompt candidates, such as Evidence-II and RolePlay-III. This overlap suggests that certain persuasive prompt strategies possess general utility, highlighting the promise of Whisper as a model-agnostic approach for efficient reasoning in black-box LRMs.
Methods MATH-500 GPQA-Diamond CommonsenseQA
Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio
DeepSeek-R1-Distill-Qwen-14B
Original 92.9 3605 100% 55.2 6799 100% 82.0 561 100%
Whisper 92.5 2521 69.9% 55.4 4255 62.6% 82.3 326 58.1%
Qwen3-14B
Original 94.5 4398 100% 66.2 7121 100% 83.8 797 100%
Whisper 95.2 2176 49.5% 66.1 3118 43.8% 84.0 328 41.2%
Table 2: Out-of-domain results of Whisper. We report both the in-domain and out-of-domain results of our best-performing prompt candidates for comparison.
Domain Generalizability
To assess the generalizability of Whisper beyond mathematical reasoning, we include two additional benchmarks in our analysis: 1) GPQA gpqa, a PhD-level science benchmark consisting of high-quality questions spanning physics, chemistry, and biology subdomains. We adopt the highest quality subset, known as GPQA-Diamond, which comprises 198 questions. 2) CommonsenseQA commonsenseqa, a widely used multiple-choice question answering dataset requiring diverse commonsense knowledge to predict correct answers. We evaluate prompts identical to those in our main experiments, which were optimized on our PDSet from the mathematical domain. As shown in Table 2, Whisper effectively reduces $\bf{37\%}$$\bf{56\%}$ of token usage on GPQA-Diamond, and achieves $\sim$$\bf{2\times}$ token reduction on CommonsenseQA. These results highlight the robust generalizability of Whisper across diverse data domains.
Figure 6: Sensitivity of different models to Whisper. Solid lines represent mean results; black diamonds indicate outliers. Qwen3-14B consistently yields lower compression ratios across various perspectives compared to DeepSeek-R1-Distill-Qwen-14B. Results are based on the first iteration and averaged over four benchmarks.
@@ 6.3 Sensitivity Analysis
Figure 6 compares the sensitivity of different model series to Whisper. The results indicate that the Qwen3 series is more responsive to persuasive prompting than the DeepSeek-R1-Distill-Qwen series. In particular, Qwen3-14B consistently achieves lower compression ratios across various persuasive perspectives compared to DeepSeek-R1-Distill-Qwen-14B, with average improvements ranging from $4\%$ to $12\%$. This trend is consistent with the findings in Table 1, where Qwen3-14B attains an average compression ratio of $63\%$, whereas DeepSeek-R1-Distill-Qwen-14B achieves $78\%$.
@@ 6.4 Impact of Refinement Iteration
We further examine the impact of iterative refinement on the performance of Whisper. As illustrated in Figure 7, successive refinement rounds improve the compression ratio of the best-performing prompt candidates. Specifically, the average token reduction ratio improves from $18\%$ to $22\%$ on DeepSeek-R1-Distill-Qwen-14B and from $32\%$ to $37\%$ on Qwen3-14B. Importantly, these refinements do not compromise the original reasoning accuracy, indicating the robustness of Whisper across multiple iterations. No further gains are observed beyond three iterations. Accordingly, we adopt three refinement rounds in our experiments.
@@ 7 Conclusion
This work presents Whisper, an iterative refinement framework that generates high-quality persuasive prompts from diverse perspectives to elicit concise LRM responses. Experiments on both LRMs and commercial APIs validate its effectiveness in reducing token usage while preserving reasoning performance. Further analysis highlights the value of different perspectives, as well as the generalizability of Whisper across diverse domains, model scales, and families. We hope this study offers new insights into efficient reasoning and underscores the potential of persuasive prompting as a practical black-box strategy for improving LRM efficiency.
@@
Due to computational constraints, we did not conduct experiments on larger LRMs such as Qwen3-235B-A22B. We believe that Whisper could retain its effectiveness in enhancing the reasoning efficiency of such models. The open-source LRMs evaluated in this study are primarily from the Qwen3 and DeepSeek-R1-Distill model series. Future work will extend our investigation to a broader range of models, including the gpt-oss series oss. Nonetheless, the strong performance of Whisper on closed-source APIs supports its generalizability across diverse model families. We plan to explore these aspects further, as we anticipate that Whisper ’s potential can be more fully realized in these settings.
@@
As stated in Section 3, our work focuses on promoting concise and efficient reasoning in reasoning models, with a clear commitment to ensuring safe and ethical usage. While our method involves techniques that may bear surface resemblance to adversarial prompting Zeng2024:persuasive; shen2024anything, particularly in their potential to influence model behavior, it fundamentally diverges from conventional jailbreak-style adversarial attacks. Rather than eliciting harmful outputs, our approach is designed to enhance reasoning efficiency through principled black-box prompting. We advocate for transparent research to promote the responsible evolution of LLM technologies.
@@
Appendix A Details of Whisper
In this section, we provide details of our proposed Whisper framework, including formal definitions of persuasive perspectives (§A.1), detailed instructions for prompt creation (§A.2), and representative examples of prompt candidates (§A.3).
@@ A.1 Definitions
We present detailed definitions of our adopted persuasive perspectives as follows:
  • Emotional Appeal. A persuasive technique that seeks to influence audiences by eliciting emotions, such as sadness, hope, or empathy, rather than relying exclusively on logic or factual evidence petty2003emotional.
  • Threat. A coercive form of persuasion that leverages explicit or implicit negative consequences, such as fear of punishment or loss, to compel behavioral or attitudinal compliance johannesen1989threat.
  • Evidence-based Persuasion. A rational appeal that utilizes credible sources, such as statistics, empirical findings, and expert testimony, to substantiate claims and influence beliefs, attitudes, or actions o2016evidence.
  • Role-Playing. A strategy involving the enactment of scenarios or personas to foster empathy, encourage perspectives, and increase the likelihood of attitude or behavior change Chen2024:rolesurvey.
Due to computational constraints, this study investigates a limited set of persuasive perspectives. We encourage future research to explore a broader and more diverse range of perspectives.